OpenStudy (anonymous):

I am taking Algebra 2 and I was wondering if anyone was good with foundations?

OpenStudy (hihi67):

Yeah, I could give them a try.

OpenStudy (anonymous):

not me lol i suck

OpenStudy (anonymous):

lol i can try help (:

OpenStudy (anonymous):

I really want to but there is no question

OpenStudy (anonymous):

1. Which ordered pair makes the open sentence 2x - y Less than Equal to10 true? (Points : 1)

OpenStudy (hihi67):

Well, there is more than one possibility.

OpenStudy (hihi67):

For example, x=5, y=2

OpenStudy (tkhunny):

Have you considered substituting the values and checking? Given (1,4), we have 2(1) - 4 = 2 - 4 = -2 < 10!!! Okay, that one works.

OpenStudy (anonymous):

(3,- 2) (4, -3) (0, -12) (7, 2)

OpenStudy (hihi67):

Okay, just plug all of those in, and see which one works.

OpenStudy (hihi67):

2 times 3 is 6. 6 -2 is 4

OpenStudy (tkhunny):

Okay, start substituting! Never "plug in". Never. Instead,, use the principle of Substitution.

OpenStudy (anonymous):

I got 3, -2?

OpenStudy (hihi67):

Yep! Great!

OpenStudy (hihi67):

Once you get it, it's not really hard, is it?